# ActionPose: Pretraining 3D Human Pose Estimation with the Dark Knowledge of Action

### v1

@@ Introduction
Experiments
Pretraining
The BABEL (Punnakkal et al. 2021) dataset is a large dataset that includes language labels describing the actions performed in mocap sequences. It provides both sequence-level labels, which describe the overall action in the sequence, and frame-level labels, which detail fine-grained actions in every frame. Each frame in the sequence can have multiple corresponding fine-grained action labels. The motion sequences in the BABEL dataset are sourced from the large mocap dataset AMASS (Mahmood et al. 2019), which uses a common parameterization framework. Instead of obtaining videos and extracting 2D skeleton sequences from the mocap dataset, we project the 3D skeletons extracted from parametric models to 2D from both a side view and a front view, assuming an orthographic camera, as done in (Zhu et al. 2023). We align the body keypoint definitions with those of Human3.6M and convert the camera coordinates to pixel coordinates using the approach outlined in (Ci et al. 2022).
@@
The overall pretraining framework consists of three layers of pretrained BERT (Devlin et al. 2018) as the text encoder and five layers of pretrained MotionBERT (Zhu et al. 2023) as the pose encoder. The network is trained on a single NVIDIA H100 GPU with a batch size of 16 and a sequence length of 243 frames and for 300 epochs.
Fine-tuning for 3D Human Pose Estimation
@@
We evaluated our methods on the public dataset Human3.6M (Ionescu et al. 2014) and MPI-INF-3DHP (Mehta et al. 2017).
@@
Human3.6M (Ionescu et al. 2014) contains 3.6 million video frames of human motion performed by professional actors in an indoor environment. Following previous works (Zhu et al. 2023; Peng, Zhou, and Mok 2024), we use subjects 1, 5, 6, 7, and 8 for fine-tuning, and subjects 9 and 11 for testing. We report results using the mean per joint position error (MPJPE), which measures the average Euclidean distance between predicted and ground-truth joint positions, typically in millimeters. Additionally, we report the Procrustes-aligned MPJPE (P-MPJPE), where MPJPE is calculated after applying Procrustes alignment to the predicted and ground-truth positions.
@@
MPI-INF-3DHP (Mehta et al. 2017) is another large-scale public dataset. This dataset contains recordings from 14 cameras capturing 8 actors performing 8 activities for the training set and 7 activities for evaluation. Following the setting of previous work (Shan et al. 2022; Mehraban, Adeli, and Taati 2024; Peng, Zhou, and Mok 2024), our evaluation metrics included the area under the curve (AUC), percentage of correct keypoints (PCK), and mean per-joint position error (MPJPE).
Performance Comparison on the Human3.6M Dataset
We evaluated ActionPose on the Human3.6M dataset (Ionescu et al. 2014) and reported MPJPE and P-MPJPE in millimeters. The 2D skeletons are extracted using Stacked Hourglass networks (Newell, Yang, and Deng 2016). The pose encoder is fine-tuned on the Human3.6M training set. As shown in Table 1, ActionPose significantly outperforms previous models based on temporal information and achieves results comparable to diffusion-based methods. Although the diffusion-based KTPFormer (Peng, Zhou, and Mok 2024) achieves better results on the Human3.6M dataset, diffusion-based methods are time-consuming during inference, as they require progressive refinement of the results.
Among all approaches, MotionBERT (Zhu et al. 2023) is the closest to us in its pretraining and fine-tuning framework. Table 1 shows that our model outperforms it by a large margin.
Performance Comparison on the MPI-INF-3DHP Dataset
To demonstrate the generalization ability of ActionPose, we evaluated its performance on the MPI-INF-3DHP dataset. Compared to Human3.6M, the test sets of MPI-INF-3DHP have more diverse motions and camera viewpoint variation.
@@
We fine-tuned the pose encoder using ground-truth 2D poses as inputs, following the approach of previous work (Mehraban, Adeli, and Taati 2024; Peng, Zhou, and Mok 2024). As shown in Table 2, ActionPose achieves state-of-the-art results, with a PCK of 98.9%, an AUC of 87.0%, and an MPJPE of 15.5mm. Notably, these results surpass all existing state-of-the-art methods, including those based on temporal information and diffusion-based approaches.
Method MPI-INF-3DHP Human3.6M
Baseline 16.7 37.5
+ MMM 16.2 36.8
+ MAP 15.7 37.3
+ MMM + MAP 15.5 36.7
@@
Table 3: Results of ablation study of each component of ActionPose on MPI-INF-3DHP and Human3.6M in MPJPE (mm). We refer to masked motion modeling as MMM and multi-modal alignment prediction as MAP.
Ablation Study
@@
To verify the effectiveness of our proposed framework, we conducted ablation studies on the MPI-INF-3DHP dataset ($T$ = 81) using ground-truth 2D poses as inputs and on the Human3.6M dataset ($T$ = 243) using detected 2D poses. Table 3 presents the results, reported as MPJPE (mm), for each component added to our framework.
@@
For the baseline, we used the pretrained MotionBERT model fine-tuned on the MPI-INF-3DHP and Human3.6M datasets, respectively. The pretrained MotionBERT weights were obtained from the official GitHub page. On MPI-INF-3DHP, pretraining with the proposed masked time windows and masked body parts motion modeling reduced the error by 0.5mm compared to the baseline. Incorporating random masking, random noise, and multi-modal alignment prediction further reduced the error by 1.0mm. Combining these strategies resulted in a total error reduction of 1.2mm from the baseline. Similarly, on Human3.6M, pretraining with masked time windows and masked body parts motion modeling reduced the error by 0.7mm, and adding random masking, noise, and multi-modal alignment reduced the error by an additional 0.2mm. The combination of these techniques led to a total error reduction of 0.8mm from the baseline.
Qualitative Results on Global Representation Learning
@@
To verify ActionPose’s ability to learn rich semantic information about actions, we visualize the global representations of pose embeddings in Figure 5. Ideally, semantically similar actions should have closer embeddings, while different actions should be farther apart. For instance, ‘throw’ and ‘pick’ both emphasize upper body movements, and thus their representations are close. In constrast, ‘jump’, which emphasizes lower body movement, is distinctly separate from them. Similarly, ‘run’ and ‘kick’ are similar from leg movements and are positioned near each other, while ‘wave’ focused on the upper body, is more distant. Additionally, ‘bend’ and ‘raise’ are closely related, whereas ‘walk’ is farther from these actions.
@@
As shown in Figure 5, ActionPose effectively groups semantically similar actions closer together while pushing semantically different actions farther apart. In contrast, MotionBERT’s global representation lacks a clear pattern, with significant overlap between both similar and distinct actions. Please note that there is some diversity within the same action due to variations in concurrent actions and different forms and styles of the actions. For example, “kicking while jumping” differs from “kicking while standing,” and the single action label with the verb ‘kick’ include motion sequences for both “kicking a soccer ball” or “kicking with the right leg back and up”.
@@
(b) ActionPose
Figure 5: The t-SNE visualization of global representations of motion embeddings on the BABEL dataset: on the left (a) are embeddings obtained through random joint-level and frame-level masking, as proposed by MotionBERT (Zhu et al. 2023), with average pooling of all pose features in the sequence; on the right (b) are embeddings obtained through ActionPose.
Discussion
Learning sufficient action semantics during text embedding alignment with motion representation requires a large dataset. When the action classes in the pretraining dataset differ from those in the target dataset, the benefits of fusing action knowledge through pretraining may diminish. Despite this limitation, ActionPose can be extended to other human motion-related tasks, such as motion prediction and completion. For motion prediction, the model can use historical motion to predict intentions, thereby guiding future movements. Similarly, for motion completion, the model can learn transitions between starting and ending poses, constrained by the corresponding actions. The framework’s ability to learn rich semantic information about actions also has potential benefits for downstream tasks such as motion captioning and generation from text descriptions. These applications are reserved for future work.
Conclusion
@@
Appendix
Additional Architecture Details
Figure 6: The architecture of pose embedding pooling layers (a) and text embedding pooling layers (b), where J represents the number of joints.
In this supplementary material, we provide a detailed illustration of the architecture for the Text Embedding and Pose Embedding Pooling Layers. As shown in Figure 6, all linear layers are activated by GELUs (Hendrycks and Gimpel 2023), followed by a layer normalization layer and/or a dropout layer. For the Pose Embedding Pooling layers, due to the extra joint dimension, we first reshape the input in the middle, then apply a weighted summation over the last dimension.
Experimental Details
Configuration
The network is implemented in PyTorch and trained on a single NVIDIA H100 GPU for 300 epochs using an AdamW optimizer, with a learning rate of 0.0005 and a batch size of 16.
Pretraining
For the pretraining datasets BABEL (Punnakkal et al. 2021) and AMASS (Mahmood et al. 2019), we first render the SMPL+H (Romero, Tzionas, and Black 2017) parametric model, then extract the 3D keypoints using a regression matrix (Bogo et al. 2016). The original AMASS dataset, which has a frequency of 120 Hz, is resampled to 30 Hz. In 50% of the random masking instances, 5% of the joints are masked at the joint level, and 15% of the frames are masked at the frame level. In 25% of the time-window masking instances, the size of the masking window is randomized between $T_{1}=30$ and $T_{2}=80$.
During pretraining, for the contrastive loss $\mathcal{L}_{con}$, the temperature is set to 0.1, and each pair of pose and text embeddings is compared against 16 negative samples.
Qualitative Results
Figure 7: Qualitative results of ActionPose compared to the baseline model MotionBERT. The left column shows the results from ActionPose, the middle column shows the results from MotionBERT for the same frames, and the right column displays the corresponding original 2D pose inputs.
Performance Comparison on the Human3.6M Dataset
In this section, we present qualitative results of our fine-tuned pose encoder on the Human3.6M dataset (Ionescu et al. 2014), comparing our results with the baseline model, MotionBERT (Zhu et al. 2023). As shown in Figure 7, ActionPose demonstrates robustness in cases of occluded 2D poses (first and third rows) and produces more accurate and natural-looking poses when 2D inputs are clear (second row).
Qualitative Results on the Pose Encoder’s Semantic Embedding
In this section, we present additional qualitative results to assess the semantic embedding capability of the fine-tuned pose encoder on the Human3.6M dataset (Ionescu et al. 2014). Specifically, we compare the similarity scores between pose embeddings and text embeddings of two selected action labels, followed by applying the SoftMax function to these scores. As shown in Figure 8, the pose encoder can distinguish not only between semantically similar actions (e.g., bend down and kneel) and semantically related actions (e.g., walk and stand), but also semantically inverse actions (e.g., sit down vs. get up from a chair, walk forward vs. walk backward).
Figure 8: Qualitative Results on the Pose Encoder’s Semantic Embedding. Each clip is manually selected from the Human3.6M dataset: (a) and (c) are from the activity ’Walking Dog’, while (b) is from ’Sitting Down’.

### v2

@@ 3.3 Fine-tuning
Method PCK $\uparrow$ AUC $\uparrow$ P1 $\downarrow$
P-STMO [42] 97.9 75.8 32.2
GLA-GCN [54] 98.5 79.1 27.7
STCFormer [47] 98.7 83.9 23.1
DiffPose [13] †* 98.0 75.9 29.1
D3DP [43] †* 97.7 78.2 29.7
KTPFormer [36] †* 98.9 85.9 16.7
MotionAGFormer-L [30] 98.2 85.3 16.2
TCPFormer [24] 99.0 87.7 15.0
RePose [46] 98.3 86.7 15.5
LangPose (Ours) $\diamond$ 98.9 87.0 15.5
4 Experiments
4.1 Pretraining
The BABEL [40] dataset is a large dataset that includes language labels describing the actions performed in mocap sequences. It provides both sequence-level labels, which describe the overall action in the sequence, and frame-level labels, which detail fine-grained actions in every frame. Each frame in the sequence can have multiple corresponding fine-grained action labels. The motion sequences in the BABEL dataset are sourced from the large mocap dataset AMASS [28], which uses a common parameterization framework. Instead of obtaining videos and extracting 2D skeleton sequences from the mocap dataset, we project the 3D skeletons extracted from parametric models to 2D from both a side view and a front view, assuming an orthographic camera, as done in [62]. We align the body keypoint definitions with those of Human3.6M and convert the camera coordinates to pixel coordinates using the approach outlined in [7].
@@ 4.1 Pretraining
The overall pretraining framework consists of three layers of pretrained BERT [9] as the text encoder and five layers of pretrained MotionBERT [62] as the pose encoder. The network is trained on a single NVIDIA H100 GPU with a batch size of 16 and a sequence length of 243 frames and for 300 epochs using an AdamW optimizer. The loss weights are set to $\lambda_{con}=0.5$, $\lambda_{3D}=1$, and $\lambda_{v}=20$, and the model is trained for 300 epochs. Please refer to the Supplementary Material for more implementation details.
4.2 Fine-tuning for 3D human pose estimation
@@
Datasets and evaluation metrics. We evaluated our methods on the public dataset Human3.6M [15] and MPI-INF-3DHP [31].
@@
Human3.6M [15] contains 3.6 million video frames of human motion performed by professional actors in an indoor environment. Following previous works [62, 36], we use subjects 1, 5, 6, 7, and 8 for fine-tuning, and subjects 9 and 11 for testing. We report results using the mean per joint position error (MPJPE), which measures the average Euclidean distance between predicted and ground-truth joint positions, typically in millimeters. We also report the Procrustes-aligned MPJPE (P-MPJPE), where MPJPE is calculated after applying Procrustes alignment to the predicted and ground-truth positions. Futhermore, we observe that three actions are consistently the most challenging to predict across state-of-the-art methods as shown in Tab. 1: Sitting (involving external objects), Sitting Down (prone to self-occlusions), and Taking Photo (less repeatable motion). We also find that these actions are the most heavily occluded, according to a rough estimation of their visibility [^1] See the Supplementary Material for details of the visibility estimation.. We denote the average MPJPE over these three actions as H-MPJPE. For simplicity, we refer to MPJPE, P-MPJPE, and H-MPJPE as P1, P2, and P3, respectively.
@@
MPI-INF-3DHP [31] is another large-scale public dataset. This dataset contains recordings from 14 cameras capturing 8 actors performing 8 activities for the training set and 7 activities for evaluation. Following the setting of previous work [42, 30, 36], our evaluation metrics included the area under the curve (AUC), percentage of correct keypoints (PCK), and mean per-joint position error (MPJPE).
Performance comparison on the Human3.6M dataset. We evaluated LangPose on the Human3.6M dataset [15], and reported P1, P2, and P3 errors in millimeters. The 2D skeletons are extracted using the Stacked Hourglass network [34], and the pose encoder is fine-tuned on the Human3.6M training set. As shown in Tab. 1, LangPose achieves competitive performance compared to prior temporal models, with particularly large improvements on the three most challenging actions. This highlights the effectiveness of incorporating semantic information when spatial and temporal consistency alone is insufficient. Moreover, LangPose achieves performance comparable to diffusion-based methods, while being more efficient during inference, as diffusion models rely on iterative refinement.
Among all baselines, MotionBERT [62] is most closely related to our framework in terms of its pretraining and fine-tuning strategy. As shown in Tab. 1, our model outperforms MotionBERT by a large margin, especially on the hardest actions, demonstrating the benefits of integrating semantic information during pretraining. It is worth noting that the motion sequences from BABEL [40] originate from the AMASS dataset [28], which is already included in the pretraining data of MotionBERT. Although our pretraining includes additional action labels to facilitate semantic alignment, no extra motion sequences are used beyond what is available to MotionBERT.
Performance comparison on the MPI-INF-3DHP dataset. To evaluate the generalization capability of LangPose, we assessed its performance on the MPI-INF-3DHP dataset. Compared to Human3.6M, the test set of MPI-INF-3DHP features more diverse motion types and greater variation in camera viewpoints. For instance, MPI-INF-3DHP includes sports-related poses that are absent from Human3.6M. Additionally, approximately 67% of the poses in Human3.6M fall under the “stand/pose” category, while only 25% correspond to “sit” and 8% to “crouch”—the latter being more complex and challenging to model [31].
@@
Following prior works [30, 36], we fine-tuned the pose encoder using ground-truth 2D poses as input. As shown in Tab. 2, LangPose achieves SOTA level performance, obtaining a PCK of 98.9%, an AUC of 87.0%, and an MPJPE of 15.5 mm. These results surpass most existing methods, including both temporal-based and diffusion-based approaches, highlighting the strong generalization ability of Pose.L
@@
Table 3: Ablation study of LangPose on MPI-INF-3DHP and Human3.6M datasets. Results are reported using P1 and P3 metrics (in mm). We refer to masked motion modeling as MMM and multi-modal alignment prediction as MAP.
Configuration Modules enabled MPI (P1) H36M (P1) H36M (P3)
MMM MAP
Baseline 16.7 37.5 49.0
Baseline + MMM 16.2 36.8 47.6
Baseline + MAP 15.7 37.3 47.9
MMM + MAP (ours) 15.5 36.7 47.0
Table 4: Ablation study comparing the proposed modified InfoNCE loss with the original InfoNCE formulation on MPI-INF-3DHP (abbrev. MPI) and Human3.6M (abbrev. H36M). MMM denotes masked motion modeling. Results are reported using P1 and P3 metrics ($m$).
Method (Loss) MPI (P1) H36M (P1) H36M (P3) Baseline (w/o InfoNCE) 16.7 37.5 49.0 Baseline + original InfoNCE + MMM 16.3 37.5 48.5 Baseline + MMM + Modified InfoNCE (ours) 15.5 36.7 47.0
@@
Ablation study. To verify the effectiveness of our proposed framework, we conducted ablation studies on the MPI-INF-3DHP dataset ($T$ = 81) using ground-truth 2D poses as inputs and on the Human3.6M dataset ($T$ = 243) using detected 2D poses. Table 3 presents the results for each component added to our framework, reported as P1 for the Human3.6M dataset and as both P1 and P3 for the MPI-INF-3DHP dataset (in mm).
@@
For the baseline, we used the pretrained MotionBERT model fine-tuned on the MPI-INF-3DHP and Human3.6M datasets, respectively. The pretrained MotionBERT weights were obtained from the official GitHub page. On MPI-INF-3DHP, pretraining with the proposed masked time windows and masked body parts motion modeling (MMM) reduced the error by 0.5mm compared to the baseline. Incorporating random frame and joint-level masking, random noise, and multi-modal alignment prediction further reduced the error by 1.0mm (MAP). Combining these strategies (MMM + MAP) resulted in a total error reduction of 1.2mm from the baseline. Similarly, on Human3.6M, pretraining with masked time windows and masked body parts motion modeling reduced the error by 0.7mm, and adding random frame and joint-level masking, noise, and multi-modal alignment reduced the error by an additional 0.2mm. The combination of these techniques led to a total error reduction of 0.8mm from the baseline. P3 is additionally reported to highlight the performance of the proposed modules on the most difficult actions in Human3.6M, where the joint contribution of masked motion modeling and semantic alignment proves especially beneficial.
To further validate the effectiveness of our proposed loss design, we conduct an additional ablation study by comparing the modified InfoNCE loss with the original InfoNCE formulation. Specifically, in the modified InfoNCE loss, the cross-entropy term is replaced with a KL-divergence objective, and a focal loss component is incorporated to better handle hard examples. As shown in  Tab. 4, this modification leads to consistent performance improvements across both benchmarks. Moreover, we observe that the original InfoNCE loss tends to suffer from overfitting during training, whereas our modified InfoNCE loss successfully mitigates this issue, providing a stronger optimization signal and yielding more robust representations for downstream 3D human pose estimation.
4.3 Qualitative results on global representation learning
@@
To verify LangPose’s ability to learn rich semantic information about actions, we visualize the global representations of pose embeddings in Fig. 4. Ideally, semantically similar actions should have closer embeddings, while different actions should be farther apart. For instance, ‘throw’ and ‘pick’ both emphasize upper body movements, and thus their representations are close. In constrast, ‘jump’, which emphasizes lower body movement, is distinctly separate from them. Similarly, ‘run’ and ‘kick’ are similar from leg movements and are positioned near each other, while ‘wave’ focused on the upper body, is more distant. Additionally, ‘bend’ and ‘raise’ are closely related, whereas ‘walk’ is farther from these actions.
@@
As shown in Fig. 4, LangPose effectively groups semantically similar actions closer together while pushing semantically different actions farther apart. In contrast, MotionBERT’s global representation lacks a clear pattern, with significant overlap between both similar and distinct actions. Please note that there is some diversity within the same action due to variations in concurrent actions and different forms and styles of the actions. For example, “kicking while jumping” differs from “kicking while standing,” and the single action label with the verb ‘kick’ include motion sequences for both “kicking a soccer ball” or “kicking with the right leg back and up”.
@@
(b) LangPose
Figure 4: The t-SNE visualization of global representations of motion embeddings on the BABEL dataset: on the left (a) are embeddings obtained through random joint-level and frame-level masking, as proposed by MotionBERT [62], with average pooling of all pose features in the sequence; on the right (b) are embeddings obtained through LangPose.
5 Discussion
Learning sufficient action semantics during text embed- ding alignment with motion representation requires a large dataset. When the action classes in the pretraining dataset differ from those in the target dataset, the benefits of fus- ing action knowledge through pretraining may diminish [^2] See the Supplementary Material for more discussion.. Despite this limitation, beyond 3D human pose estimation, LangPose can be extended to other human motion-related tasks, such as motion prediction and completion. For motion prediction, the model can use historical motion to predict intentions, thereby guiding future movements. Similarly, for motion completion, the model can learn transitions between starting and ending poses, constrained by the corresponding actions. The framework’s ability to learn rich semantic information about actions also has potential benefits for downstream tasks such as motion captioning from noisy inputs and generation from text descriptions. These applications are reserved for future work.
6 Conclusion
@@
Figure 5: Qualitative results of LangPose compared to the baseline model MotionBERT. The left column shows the results from LangPose, the middle column shows the results from MotionBERT for the same frames, and the right column displays the corresponding original 2D pose inputs.